OpenStudy (anonymous):

I know this is easy, but I forgot. :P If b^2 + d^2 = 21 and (b+d)^2 = 25, then what is bd?

jimthompson5910 (jim_thompson5910):

(b+d)^2 = 25 b^2 + 2bd + d^2 = 25 2bd + (b^2 + d^2) = 25 2bd = 25 - (b^2 + d^2) 2bd = 25 - (21) I'm sure you see what to do from here

OpenStudy (anonymous):

Yeah I do, thanks!

jimthompson5910 (jim_thompson5910):

you're welcome